# Parallelizing MCMC Across the Sequence Length

### v1

@@ A.1 HMC algorithm with parallel leapfrog
A.2 Block quasi-DEER for HMC with Parallel Leapfrog
[…]
@@ A.2 Block quasi-DEER for HMC with Parallel Leapfrog
$\boldsymbol{J}_{BQ}(\mathbf{s}_{t-1})=\begin{bmatrix}\mathrm{diag}(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{x}_{t-1}})&\mathrm{diag}(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{v}_{t-1}})\\ \mathrm{diag}(\frac{\partial\mathbf{v}_{t}}{\partial\mathbf{x}_{t-1}})&\mathrm{diag}(\frac{\partial\mathbf{v}_{t}}{\partial\mathbf{v}_{t-1}}).\end{bmatrix}$ (15)
@@
[…]
@@ Appendix B Additional Experiment Details
In our applications we use a combination of absolute and relative tolerance [Hairer et al., 1993] as the stopping criterion for parallel Newton’s method. In all experiments, we set the absolute tolerance to $10^{-4}$ and relative tolerance to $10^{-3}$ unless otherwise noted.
@@ B.1 Parallel MALA
We use Bayesian logistic regression (BLR) on the German Credit Dataset (with whitened covariates) [Dua et al., 2017, TFDS, ] as our testbed for this set of experiments. Specifically, we will use quasi-DEER-based Parallel MALA to generate $B$ independent chains of $L$ samples each from the posterior distribution of the BLR model with a $\mathcal{N}(\mathbf{0},\boldsymbol{I})$ prior on the covariates. We vary the batch size $B\in\{1,2,4,8,16,32\}$ and the sequence length $L\in\{1000,2000,4000,8000,16000,32000,64000\}$. To solidify confidence in our results, we will repeat all settings across 20 random seeds. We repeat the corresponding settings for the baseline sequential MALA algorithm. To ensure fairness, we implement the sequential procedure using the JIT-compilable jax.lax.scan module.
@@ B.1.1 Measuring Sample Quality via Maximum Mean Discrepancy (MMD)
To assess sample quality, we use Maximum Mean Discrepancy [Gretton et al., 2012]. We do not use traditional MCMC metrics like effective sample size (ESS) or $\hat{R}$ because such metrics mainly target the autocorrelatedness of traditional sequential MCMC algorithms. In contrast, our parallel MALA algorithm by nature does not proceed sequentially sample-by-sample, and the autocorrelation of its generated samples are inherently limited by the autocorrelation of the base sequential MALA algorithm.
@@
[…]
@@
However, we do not have access to the true target posterior distribution $p_{2}$. Thus, we approximate the “ground truth” $p_{2}$ via a large set of gold standard MCMC samples generated using the state-of-the-art No U-Turn Sampler with Dual-Averaging Stepsize Adaptation introduced by [Hoffman et al., 2014] and implemented in TensorFlow Probability [Dillon et al., 2017]. Specifically, following standard best practices, we aim for a target acceptance probability of $75\%$ with initial states being drawn from the $\mathcal{N}(\mathbf{0},\boldsymbol{I})$ prior. To ensure high quality, we perform $50,000$ burn-in steps before collecting $100,000$ samples as our gold standard approximation of the true target posterior.
@@
with hyperparameter $\sigma$ to be discussed in the next paragraph. Let $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ refer to the samples outputted by sequential or parallel MALA, with $N=B\times L$ and let $\{\boldsymbol{y}_{j}\}_{j=1}^{M}$ represent the gold standard samples. In such a setting, we can empirically and unbiasedly estimate the MMD of our samples with respect to the true target posterior as follows (see [Gretton et al., 2012]):
$\widehat{\text{MMD}}(\mathcal{F}_{k},p_{1},p_{2})=\frac{1}{N(N-1)}\sum_{i=1}^{N}\sum_{j\neq i}^{N}k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})+\frac{1}{M(M-1)}\sum_{i=1}^{M}\sum_{j\neq i}^{M}k(\boldsymbol{y}_{i},\boldsymbol{y}_{j})-\frac{2}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}k(\boldsymbol{x}_{i},\boldsymbol{y}_{j}).$
@@ B.2 Gibbs sampling experiment
For each chain, we first simulated a draw from the prior distribution and then ran one sequential sampling step targeting the posterior. The result of this sequential sampling step was set to be the initial state $\mathbf{s}_{0}$ for each chain. For Newton’s method, the guess of the state sequence at iteration 0 was set to the initial state such that $\mathbf{s}_{t}^{(0)}=\mathbf{s}_{0}$ for all $t=1,\ldots,T$, as in the MALA experiment.
When using quasi-DEER for this problem, we used a 3-sample stochastic estimate of the diagonal. We additionally applied a tuned diagonal preconditioner to the stochastic quasi-DEER Jacobian estimates.
To compare the timing of sequential and parallel sampling, we ran chains of batch size $B\in\{1,2,4,8,16,$ 32,64}32,64\} and length $L\in\{8000,16000,32000,64000,128000\}$. For all batch sizes except for $B=64$, we additionally simulated chains of length $L\in\{256000,512000,1024000\}$. On a CPU, we ran sequential sampling for all batch sizes and for lengths up to $L=32000$; for values of $L>32000$ we extrapolated the time for sequential sampling (i.e. for $L=64000$ we doubled the amount of time it took to simulate $L=32000$ samples). We repeated each run 5 times and computed the timing in a similar manner as our MALA experiments. Namely, we computed the average time to run the JIT-compiled code across 3 timing runs after performing 2 warm-up runs of each algorithm.
@@ B.3 Sequential HMC with parallel leapfrog experimental details
We also applied sequential HMC with parallel leapfrog to the Bayesian logistic regression model of the German-Credit dataset Dua et al. [2017], TFDS with whitened covariates. We focused on sampling 4 batched chains and swept across the number of leapfrog steps $L=\{4,8,12,16,20,24,32,40,64,96\}$ and step sizes $\epsilon=\{0.005,0.0075,0.01,0.02,0.03,0.04,0.05,0.075,0.1\}$. Step sizes at $0.1$ and above resulted in unstable integration and acceptance rates near zero. For each combination of $L$ and $\epsilon$, we estimated the time to simulate 2000 samples from HMC using either sequential or parallel leapfrog integration using a JIT-compiled function. We repeated each step size and number of leapfrog steps across 10 random seeds. In all cases, we only used the absolute tolerance. For each run, we computed the average effective sample size across parameters. We then reported the average effective sample size per second. A relative comparison of the average ESS/s across runs between HMC with parallel or sequential leapfrog is reported in Figure 5B. Additionally, the maximum ESS/s across settings of $L$ and $\epsilon$ for HMC with parallel or sequential leapfrog is shown in Figure 5C.
@@
We next explored using sequential HMC with parallel leapfrog for an item-repsonse theory model Gelman and Hill [2007], TFDS . The target distribution in this case has $D=501$ dimensions and the dataset has approximately 3030K datapoints, making this a substantially larger problem. We compared the time to run HMC with either sequential or parallel leapfrog across 1 or 2 chains and varying step sizes and number of leapfrog steps. For each setting we ran 10 repeats across different random number seeds. We estimated the time to simulate 1000 samples from the posterior distribution (Figure 19). Here, we used a pre-tuned diagonal mass matrix, which empirically we found helped improve the convergence of parallel leapfrog integration. We found that when simulating 1 chain the HMC with parallel leapfrog can achieve speed-ups over HMC with sequential leapfrog. However, at two chains the sequential leapfrog was faster.
@@ B.3.1 Step-size for HMC with parallel leapfrog
Tuning the step size and number of leapfrog in HMC is important for optimal exploration of the target distribution. The standard analyses for determining the optimal step size and acceptance rate typically assume a linear time and computational cost in the number of leapfrog steps [Neal et al., 2011, Beskos et al., 2013]. However, parallelizing the leapfrog integration can enable sublinear cost in the number of steps, violating this assumption. A second compounding factor is that the number of parallel Newton iterations until convergence of the parallel integration can vary as a function of the step size, as we saw in our results. New analysis investigating the appropriate optimal step sizes and acceptance rates need to be devised under these conditions. We hypothesize that in some cases, these conditions will favor the use of smaller step sizes and more leapfrog steps, thereby increasing the optimal acceptance rate. We currently leave a formal analysis of this to future work.
@@ B.4 Parallel HMC across the sequence
We parallelized HMC across the sequence for both the Bayesian logistic regression example and for the Rosenbrock distribution.
For parallel HMC on Bayesian logistic regression we used memory-efficient quasi-DEER to parallelize the algorithm. We tuned the step size, number of leapfrog steps, and hyperparameters of the algorithm to consistently achieve fast convergence with high numbers of samples. The resulting parameters were $L=4$ leapfrog steps and step size $\epsilon=0.04$. We simulated five sequential samples before setting the initial state for parallel HMC. In this case, we simulated batches of $4$ chains of length $1000$. The initial guess of the states for parallel Newton’s method was set to the initial state for all time points. When using memory-efficient quasi-DEER, we clipped the magnitude of the diagonal Jacobian to be less than 1.0 and we used a damping factor of 0.4 [Gonzalez et al., 2024]. We repeated the sampler across 10 different seeds. The sampler typically converged in around 20 iterations and the acceptance ratio was around $0.89\%$. We computed the ESS/s in the same manner as for sequential HMC with sequential or parallel leapfrog in the previous subsection.
In Figure 1, we parallelized HMC sampling of the Rosenbrock “banana” distribution implemented in TFDS . We parallelized a chain of 100K samples where each HMC consisted of $L=8$ leapfrog steps with step size $\epsilon=0.5$. This resulted in an acceptance rate of approximately $97.6\%$. For this 2D distribution, we used the full DEER Jacobian scaled by a damping factor of $0.5$ [Gonzalez et al., 2024] and clipped to be within [1,1][-1,1]. For this motivating example, we ran this with a single random seed and did not examine timing or convergence profiles across runs.

### v2

@@ A.2 HMC algorithm with parallel leapfrog
A.3 Block quasi-DEER for HMC with Parallel Leapfrog
Here we discuss parallelizing leapfrog with block quasi-DEER in additional detail. Recall that for HMC with parallel leapfrog, the state of the function being parallelized is the concatenation of the position and momentum states $\mathbf{s}_{t}=[\mathbf{x}_{t},\mathbf{v}_{t}]$ with $\mathbf{s}_{t}\in\mathbb{R}^{2D}$. The Jacobian of the leapfrog step used in HMC with parallel leapfrog integration has a block structure:
@@ A.3 Block quasi-DEER for HMC with Parallel Leapfrog
$\boldsymbol{J}_{BQ}(\mathbf{s}_{t-1})=\begin{bmatrix}\mathrm{diag}(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{x}_{t-1}})&\mathrm{diag}(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{v}_{t-1}})\\ \mathrm{diag}(\frac{\partial\mathbf{v}_{t}}{\partial\mathbf{x}_{t-1}})&\mathrm{diag}(\frac{\partial\mathbf{v}_{t}}{\partial\mathbf{v}_{t-1}})\end{bmatrix}$. (14)
For parallelizing leapfrog integration in HMC, this approximate Jacobian is
@@
$\mathrm{diag}(\nabla_{\mathbf{x}}^{2}\log p(\mathbf{x}_{t-1}+\epsilon\mathbf{v}_{t-1}))=\mathbb{E}_{\mathbf{z}\sim\operatorname{Rad}}[\mathbf{z}\odot\big(\nabla_{\mathbf{x}}^{2}\log p(\mathbf{x}_{t-1}+\epsilon\mathbf{v}_{t-1})\,\mathbf{z}\big)].$ (16)
@@ Appendix B Additional Experiment Details
In our applications we use a combination of absolute and relative tolerance [80] as the stopping criterion for parallel Newton’s method. In all experiments, we set the absolute tolerance to $10^{-4}$ and relative tolerance to $10^{-3}$ unless otherwise noted.
@@ B.1 Parallel MALA
We use Bayesian logistic regression (BLR) on the German Credit Dataset (with whitened covariates) [60, 61] as our testbed for this set of experiments. Specifically, we will use quasi-DEER-based Parallel MALA to generate $B$ independent chains of $L$ samples each from the posterior distribution of the BLR model with a $\mathcal{N}(\mathbf{0},\boldsymbol{I})$ prior on the covariates. We vary the batch size $B\in\{1,2,4,8,16,32\}$ and the sequence length $L\in\{1000,2000,4000,8000,16000,32000,64000\}$. To solidify confidence in our results, we will repeat all settings across 20 random seeds. We repeat the corresponding settings for the baseline sequential MALA algorithm. To ensure fairness, we implement the sequential procedure using the JIT-compilable jax.lax.scan module.
@@ B.1.1 Measuring Sample Quality via Maximum Mean Discrepancy (MMD)
To assess sample quality, we use Maximum Mean Discrepancy [62]. We do not use traditional MCMC metrics like effective sample size (ESS) or $\hat{R}$ because such metrics mainly target the autocorrelatedness of traditional sequential MCMC algorithms. In contrast, our parallel MALA algorithm by nature does not proceed sequentially sample-by-sample, and the autocorrelation of its generated samples are inherently limited by the autocorrelation of the base sequential MALA algorithm.
@@
$\text{MMD}(\mathcal{F},p_{1},p_{2})=\sup_{f\in\mathcal{F}}\left(\mathbb{E}_{p_{1}}[f(X)]-\mathbb{E}_{p_{2}}[f(Y)]\right),$
@@
However, we do not have access to the true target posterior distribution $p_{2}$. Thus, we approximate the “ground truth” $p_{2}$ via a large set of gold standard MCMC samples generated using the state-of-the-art No U-Turn Sampler with Dual-Averaging Stepsize Adaptation introduced by [19] and implemented in TensorFlow Probability [81]. Specifically, following standard best practices, we aim for a target acceptance probability of $75\%$ with initial states being drawn from the $\mathcal{N}(\mathbf{0},\boldsymbol{I})$ prior. To ensure high quality, we perform $50,000$ burn-in steps before collecting $100,000$ samples as our gold standard approximation of the true target posterior.
@@
with hyperparameter $\sigma$ to be discussed in the next paragraph. Let $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ refer to the samples outputted by sequential or parallel MALA, with $N=B\times L$ and let $\{\boldsymbol{y}_{j}\}_{j=1}^{M}$ represent the gold standard samples. In such a setting, we can empirically and unbiasedly estimate the MMD of our samples with respect to the true target posterior as follows (see [62]):
$\widehat{\text{MMD}}(\mathcal{F}_{k},p_{1},p_{2})$ $=\frac{1}{N(N-1)}\sum_{i=1}^{N}\sum_{j\neq i}^{N}k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$
$\quad+\frac{1}{M(M-1)}\sum_{i=1}^{M}\sum_{j\neq i}^{M}k(\boldsymbol{y}_{i},\boldsymbol{y}_{j})-\frac{2}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}k(\boldsymbol{x}_{i},\boldsymbol{y}_{j}).$
@@ B.2 Gibbs sampling experiment
For each chain, we first simulated a draw from the prior distribution and then ran one sequential sampling step targeting the posterior. The result of this sequential sampling step was set to be the initial state $\mathbf{s}_{0}$ for each chain. For Newton’s method, the guess of the state sequence at iteration 0 was set to the initial state such that $\mathbf{s}_{t}^{(0)}=\mathbf{s}_{0}$ for all $t=1,\ldots,T$, as in the MALA experiment. When using quasi-DEER for this problem, we used a 3-sample stochastic estimate of the diagonal. We additionally applied a tuned diagonal preconditioner to the stochastic quasi-DEER Jacobian estimates.
To compare the timing of sequential and parallel sampling, we ran chains of batch size $B\in\{1,2,4,8,16,$ 32,64}32,64\} and length $L\in\{8000,16000,32000,64000,128000\}$. For all batch sizes except for $B=64$, we additionally simulated chains of length $L\in\{256000,512000,1024000\}$. On a CPU, we ran sequential sampling for all batch sizes and for lengths up to $L=32000$; for values of $L>32000$ we extrapolated the time for sequential sampling (i.e. for $L=64000$ we doubled the amount of time it took to simulate $L=32000$ samples). We repeated each run 5 times and computed the timing in a similar manner as our MALA experiments. That is, we computed the average time to run the JIT-compiled code across 3 timing runs after performing 2 warm-up runs of each algorithm.
@@ B.3 Sequential HMC with parallel leapfrog experimental details
We also applied sequential HMC with parallel leapfrog to the Bayesian logistic regression model of the German-Credit dataset [60, 61] with whitened covariates. We focused on sampling 4 batched chains and swept across the number of leapfrog steps $L=\{4,8,12,16,20,24,32,40,64,96\}$ and step sizes $\epsilon=\{0.005,0.0075,0.01,0.02,0.03,0.04,0.05,0.075,0.1\}$. Step sizes at $0.1$ and above resulted in unstable integration and acceptance rates near zero. For each combination of $L$ and $\epsilon$, we estimated the time to simulate 2000 samples from HMC using either sequential or parallel leapfrog integration using a JIT-compiled function. We repeated each step size and number of leapfrog steps across 10 random seeds. In all cases, we only used the absolute tolerance. For each run, we computed the average effective sample size across parameters. We then reported the average effective sample size per second. A relative comparison of the average ESS/s across runs between HMC with parallel or sequential leapfrog is reported in Figure 5B. Additionally, the maximum ESS/s across settings of $L$ and $\epsilon$ for HMC with parallel or sequential leapfrog is shown in Figure 5C.
@@
We next explored using sequential HMC with parallel leapfrog for an item-repsonse theory model [64, 61]. The target distribution in this case has $D=501$ dimensions and the dataset has approximately 3030K datapoints, making this a substantially larger problem. We compared the time to run HMC with either sequential or parallel leapfrog across 1 or 2 chains and varying step sizes and number of leapfrog steps. For each setting we ran 10 repeats across different random number seeds. We estimated the time to simulate 1000 samples from the posterior distribution (Figure 19). Here, we used a pre-tuned diagonal mass matrix, which empirically we found helped improve the convergence of parallel leapfrog integration. We found that when simulating 1 chain the HMC with parallel leapfrog can achieve speed-ups over HMC with sequential leapfrog. However, at two chains the sequential leapfrog was faster.
@@ B.3.1 Step-size for HMC with parallel leapfrog
Tuning the step size and number of leapfrog in HMC is important for optimal exploration of the target distribution. The standard analyses for determining the optimal step size and acceptance rate typically assume a linear time and computational cost in the number of leapfrog steps [12, 82]. However, parallelizing the leapfrog integration can enable sublinear cost in the number of steps, violating this assumption. A second compounding factor is that the number of parallel Newton iterations until convergence of the parallel integration can vary as a function of the step size, as we saw in our results. New analysis investigating the appropriate optimal step sizes and acceptance rates need to be devised under these conditions. We hypothesize that in some cases, these conditions will favor the use of smaller step sizes and more leapfrog steps, thereby increasing the optimal acceptance rate. We currently leave a formal analysis of this to future work.
@@ B.4 Parallel HMC across the sequence
We parallelized HMC across the sequence for both the Bayesian logistic regression example and for the Rosenbrock distribution. For parallel HMC on Bayesian logistic regression we used memory-efficient quasi-DEER to parallelize the algorithm. We tuned the step size, number of leapfrog steps, and hyperparameters of the algorithm to consistently achieve fast convergence with high numbers of samples. The resulting parameters were $L=4$ leapfrog steps and step size $\epsilon=0.04$. We simulated five sequential samples before setting the initial state for parallel HMC. In this case, we simulated batches of $4$ chains of length $1000$. The initial guess of the states for parallel Newton’s method was set to the initial state for all time points. When using memory-efficient quasi-DEER, we clipped the magnitude of the diagonal Jacobian to be less than 1.0 and we used a damping factor of 0.4 [8]. We repeated the sampler across 10 different seeds. The sampler typically converged in around 20 iterations and the acceptance ratio was around $0.89\%$. We computed the ESS/s in the same manner as for sequential HMC with sequential or parallel leapfrog in the previous subsection.
In Figure 1, we parallelized HMC sampling of the Rosenbrock “banana” distribution implemented in TFDS [61]. We parallelized a chain of 100K samples where each HMC consisted of $L=8$ leapfrog steps with step size $\epsilon=0.5$. This resulted in an acceptance rate of approximately $97.6\%$. For this 2D distribution, we used the full DEER Jacobian scaled by a damping factor of $0.5$ [8] and clipped to be within [1,1][-1,1]. For this motivating example, we ran this with a single random seed and did not examine timing or convergence profiles across runs.